Working class- Mean: 12.80 Standard Dev: 2.85 N: 697
Upper class- Mean: 15.45 Standard Dev: 2.98 N: 38
1) Assuming that years of education is normally distributed to the population, what proportion of working class respondents have 12 to 16 years of education?
2) What proportion of upper-class respondents have 12 to 16 years of education?

Answer by Boreal(15235)

z=(x-mean)/sd
z=(12.00-12.80)/2.85;; This is - 0.80/2.85=-0.28
z=(16.00-12.80)/2.85;; this is 3.20/2.85= 1.12
probability in this range is 0.4789;
z=(12-15.45)/2.98;;this is -1.16
z=(16-15.45)/2.98;;this is 0.18
probability is 0.4483
The probability is heavily weighted towards the right end for the first and towards the left for the second. Another way of saying it as that most of the probability is fewer than 12 years for the first group and greater than 16 years for the second.